OpenStudy (anonymous):

What is the area of the actual living room in square feet? Round your answer to the nearest whole number. 17 ft2 41 ft2 495 ft2 103 ft2

OpenStudy (danjs):

What are the dimensions?

OpenStudy (anonymous):

I dont know what you mean

OpenStudy (danjs):

Is there a picture to go along with the problem? or anything

OpenStudy (anonymous):

yeah here is the pic of the problem

OpenStudy (danjs):

Ok So the area of a rectangle is the length times the width

OpenStudy (danjs):

Area = 5.5*3 = 16.5

OpenStudy (danjs):

and the drawing is at a scale of 1:30

OpenStudy (anonymous):

what does i mean with the scale

OpenStudy (danjs):

1 inch on the drawing is 30 inches in real life

OpenStudy (danjs):

multiply both the sides by 30 and multiply them to find the area in real life

OpenStudy (anonymous):

that would be 495

OpenStudy (danjs):

5.5 * 30 = 165 in divide by 12 to get feet = 13.75ft 3*30 = 90 in divide by 12 to get ft = 7.5 ft

OpenStudy (anonymous):

uhhhhhhh

OpenStudy (danjs):

now the area in square feet is 13.75 * 7.5 = 103.125 ft^2

OpenStudy (danjs):

so it is 103 ft^2

OpenStudy (anonymous):

oh thx

OpenStudy (danjs):

u understand?

OpenStudy (anonymous):

yes thank you